You have just purchased 200 shares of stock on margin at $60 per share. The broker lent you $5,000 to help pay for the purchase.

a. The initial margin in the account at the time you purchased the stock is 7000

b. If the share price falls to $35 per share, the remaining margin in the account is ______

c. If the maintenance margin requirement is 35%, the amount of the margin call will be _____
Answer #1

Value of Position = 200(6,000) = $12,000

Broker Amount = $5,000

Initial margin = 12,000 - 5,000 = $7,000

a.

Initial Margin 7,000/12,000 = 58.33%

b.

Remaining Margin = 35(200) - 5,000

Remaining Margin = $2,000

c.

Margin Call Price = 60(1 - 0.5833)/(1 - 0.35)

Margin Call Price = $38.46